3a over 4a-20 + 9a over 6a-30 =
show work

Answer by fractalier(6550)

Posted this one yesterday...
3a / 4a-20 + 9a / 6a-30 =
9a / 12a - 60 + 18a / 12a - 60 =
27a / 12a - 60 =
27a / 12(a - 5) =
9a / 4(a - 5)